Ask your own question, for FREE!
Probability 8 Online
OpenStudy (anonymous):

A company has a security system comprising four electronic devices (A, B, C, and D) which operate independently. Each device has a probability of failure of 0.1. The four electronic devices are arranged so that the whole system operates properly if at least one of A or B functions and at least one of C or D functions. What is the probability that the whole system functions properly?

OpenStudy (anonymous):

So there is a probability of .1+.1=.2 that A or B will fail. Likewise, there is a probability of .1+.1=.2 that C or D will fail. So there is a (1-.2)(1-.2)=.64 chance that the system works. Basically when you see or you add probability and when you see and you multiply probability. Someone with more knowledge in probability than I can expand upon this if you want to know the intricacies.

OpenStudy (ybarrap):

|dw:1380061043565:dw| Probability of failure is if A and B fail OR C and D fail Probability of A and B failing is \(0.1^2\). Probability of C and D failing is \(0.1^2\). Probability that system functions is 1 - Probability of Failure = 1 - (Probability of A and B failing OR Probability of C and D failing) = \(1-2\times.1^2=0.98\) $$ Therefore, the probability that the system functions properly is 0.98.

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!